OpenStudy (anonymous):

suppose a baseball player had 212 hits in a season. In the given probability distribution, the random variable X represents the number of hits the player obtained in a game x: 0, 1, 2, 3, 4, 5 p(x): 0.1147, 0.4977, 0.2822, 0.0681, 0.0131, 0.0242. Compute the standard deviation of the random variable x.

OpenStudy (kropot72):

Are you supposed to use a computer program of a calculator for this?

OpenStudy (anonymous):

You can. I believe there are other ways as well.

OpenStudy (kropot72):

If you want to calculate it manually I can help. However I must log out for about an hour and will help after that if necessary :)

OpenStudy (anonymous):

I believe I found out how to solve it. I had to find the mean and then I used a normal distribution table to find the answer. If it is not correct let me know. The answer I got is 0.9251

OpenStudy (kropot72):

What value did you get for the mean?

OpenStudy (kropot72):

The question does not say that the probability distribution is normal.

OpenStudy (kropot72):

@Meg1985 Are you there?

OpenStudy (anonymous):

oh I didn't catch that

OpenStudy (anonymous):

my answer is wrong then

OpenStudy (kropot72):

Just to confirm, the correct value for the mean is 1.9129. What did you get?

OpenStudy (anonymous):

used formula (0x9)+(1x9.2)+(2X10.1)+(3x11.1)...so on

OpenStudy (kropot72):

In this case the mean is \[(0\times0.1147)+(1\times 0.4977)+(2\times 0.2822)+(3\times 0.0681)+(4\times 0.131)+(5\times 0.0245)=1.9129\]

OpenStudy (kropot72):

The missing part of the summation is \[...+(5\times 0.0245)=1.9129\]

OpenStudy (kropot72):

The variance (the square of the standard deviation) for a random variable X taking the values \[x _{1}, x _{2}, x _{3},.....x _{n}\] with probabilities \[p(x _{1}),p(x _{2}),...p(x _{n})\] is given by \[Var(X)=\sum_{i=1}^{n}(x _{i}-\mu)^{2}p(x _{i})\] Do you want me to help you thru this summation?

OpenStudy (anonymous):

thank you but I got it now. Thank you so much!

OpenStudy (anonymous):

It was the formula that I was missing

OpenStudy (kropot72):

You're welcome :)